Estimate the cost (in dollars) to build a new plant in 2018 if the index for this type of equipment has increased at an average rate of 3 % per year for the past 10 years


QUESTION 4 

A small textile plant was constructed in 2008. The major equipment, costs, and factors are shown in the table below.

EquipmentReference SizeReference CostCost-Capacity FactorNew Design Size
Finishing machine150yd//min$985,0000.82200yd//min
Jet dyer200yd//min$1,230,0000.77450yd//min
Steam dyer100yd//min$820,0000.69175yd//min

Estimate the cost (in dollars) to build a new plant in 2018 if the index for this type of equipment has increased at an average rate of 3 % per year for the past 10 years. (Enter your answer as a number without the dollar $ sign.)
Answer #1

C = c (Q/q)n

C = Cost of new design equipment

c = Cost of old equipment

Q = Size of new equipment

q = size of old equipment

n = Cost capacity factor

1) c = 985000

q = 150 yd/min

Q = 200 yd / min

n = 0.82

C = 985000(200/150)0.82

C = 1247056.11$

Equipment q (yd/min) c($) n Q(yd/min) C = c (Q/q)n
FLYING MACHINE 150 985000 0.82 200 1247056.11
Jet dyer 200 1230000 0.77 450 2296601.53
Steam dyer 100 820000 0.69 175 1206452.16